# CRANE: Reasoning with constrained LLM generation

### v3

@@ Abstract
Code generation, symbolic math reasoning, and other tasks require LLMs to produce outputs that are both syntactically and semantically correct. Constrained LLM generation is a promising direction to enforce adherence to formal grammar, but prior works have empirically observed that strict enforcement of formal constraints often diminishes the reasoning capabilities of LLMs. In this work, we first provide a theoretical explanation for why constraining LLM outputs to very restrictive grammars that only allow syntactically valid final answers reduces the reasoning capabilities of the model. Second, we demonstrate that by augmenting the output grammar with carefully designed additional rules, it is always possible to preserve the reasoning capabilities of the LLM while ensuring syntactic and semantic correctness in its outputs. Building on these theoretical insights, we propose a reasoning-augmented constrained decoding algorithm, CRANE, which effectively balances the correctness of constrained generation with the flexibility of unconstrained generation. Experiments on multiple open-source LLMs and benchmarks show that CRANE significantly outperforms both state-of-the-art constrained decoding strategies and standard unconstrained decoding, showing up to 10% points accuracy improvement over baselines on challenging symbolic reasoning benchmarks GSM-symbolic and FOLIO.
@@ 2.2 Deterministic LLM Decoding
Under constrained decoding with a formal grammar $G$, the output string $\boldsymbol{y}_{G}$ is selected from the constrained output distribution and is denoted as $\boldsymbol{y}_{G}=\mathcal{L}{}_{G}(\boldsymbol{x})$. The output of $i$-th constrained autoregressive step with $G$ is $\boldsymbol{x}\cdot\boldsymbol{y}^{(i)}_{G}=\mathcal{L}{}_{G}^{(i)}(\boldsymbol{x})$ and $\boldsymbol{x}\cdot\boldsymbol{y}_{G}=\mathcal{L}{}_{G}^{(|\boldsymbol{y}|)}(\boldsymbol{x})$.
The constrained output $\boldsymbol{y}_{G}$ is always in the grammar $\boldsymbol{y}_{G}\in L(G)$ where $L(G)$ is the language defined by $G$. For sound-constrained decoding algorithms, if the unconstrained output $\boldsymbol{y}=\mathcal{L}(\boldsymbol{x})$ in the grammar $\boldsymbol{y}\in L(G)$, the constrained output remains unchanged, i.e., $\mathcal{L}(\boldsymbol{x})=\mathcal{L}{}_{G}(\boldsymbol{x})$.
@@ A Turing machine $M$ with $k$ work tapes and an output tape is a $8$-tuple
where $\Sigma$ is the finite input alphabet, $\Gamma$ is the finite tape alphabet with $\Sigma\subseteq\Gamma$, $b\in\Gamma\setminus\Sigma$ is a special blank symbol, $Q$ is a finite set of states, $q_{0}\in Q$ is the initial state, $\delta:\left(Q\setminus F\right)\times\Gamma^{k+2}\to Q\times\Gamma^{k+1}\times\{0,+1,-1\}^{k+2}$ is the transition function (where $-1,1,0$ represent moving the tape head left, right, or staying in place, respectively), and $F\subseteq Q$ is the set of halting states.
Let $\Sigma^{*}$ denote the set of all finite strings over the input alphabet $\Sigma$. Given an input string $\boldsymbol{s}\in\Sigma^{*}$, the computation of $M$on $s$ is a sequence of configurations starting from the initial configuration. Each configuration $\gamma$ is a tuple containing the current state $q\in Q$, the contents of the input tape, the $k$ work tapes, the output tape, and the current head positions of all $k+2$ tapes. For each configuration, $\gamma_{i}$ ($i\in\mathbb{N}$), the transition function $\delta$ computes the next configuration $\gamma_{i+1}$ based on the current state $q$ and the values on the $k+2$ tapes at the current head positions. It updates the head positions, writes to the output tape (possibly leaving it unchanged if no new symbol is written), and advances to the next configuration. For each $i$, computation of $\gamma_{i+1}$ from $\gamma_{i}$ defines a single step of the Turing machine.
@@ 3.1 Limitation of Constrained Decoding
Let $\boldsymbol{x}\cdot\boldsymbol{y}^{(i)}$ denote the output after the $i$-th autoregressive step of an LLM $\mathcal{L}{}$ under constrained decoding with an output grammar $G$ on input $\boldsymbol{x}$. Then, we have $\boldsymbol{x}\cdot\boldsymbol{y}^{(i)}=\mathcal{L}{}_{G}^{(i)}(\boldsymbol{x})$, and for any $i$, $\boldsymbol{y}^{(i)}$ is always a valid prefix of a string in $L(G)$, i.e., there exists a (possibly empty) string $\boldsymbol{\alpha}^{(i)}$ such that $\boldsymbol{y}^{(i)}\cdot\boldsymbol{\alpha}^{(i)}\in L(G)$. Now, for any output grammar $G_{c}$ where the output set $O=L(G_{c})$ is finite, we show that the output ($\mathcal{L}{}_{G_{c}}(\boldsymbol{x})$ for any input $\boldsymbol{x}$ of size $|\boldsymbol{x}|=n$ can be computed using constant-depth threshold circuits.
@@
For any log-precision LLM $\mathcal{L}{}$ with constant layers there exists a logspace-uniform thershold circuit $Th_{n}$ such that ($\mathcal{L}{}_{G_{c}}(\boldsymbol{x})=Th_{n}(\boldsymbol{x})$ holds for all inputs $\boldsymbol{x}$ with size $|\boldsymbol{x}|=n$ and $n\in\mathbb{N}$.
@@ 3.2 Reasoning with Augmented Grammar
At a high level, $\mathcal{L}{}_{M}$ simulates the Turing machine $M$ by computing the encoded representations $\overline{\gamma_{i}}$ of the machine’s configurations $\gamma_{i}$ at each step $i$ and storing them within the reasoning component (i.e., the string $\boldsymbol{r}$) of the output. During each autoregressive step, $\mathcal{L}_{M}$ generates the next configuration based on the transition function of $M$ and appends its encoding to the reasoning sequence. This process continues until $M$ reaches a halting state, at which point $\mathcal{L}_{M}$ produces the final output $\boldsymbol{a}$, which belongs to $L(G)$. For any given $M$, we define the rules $R_{M}$ that can parse the encodings $\overline{\gamma}$ of all possible configurations $\gamma$. This ensures that the output ($\mathcal{L}{}_{G_{a}}(\boldsymbol{x})$ represents the full reasoning-augmented sequence, i.e., $\overline{\gamma_{1}}\cdots\overline{\gamma_{t(n)}}\cdot M(\boldsymbol{x})$, where $M(\boldsymbol{x})$ is the final output of $M$ on input $\boldsymbol{x}$ of size $n$ after $t(n)$ computational steps. The encodings $\overline{\gamma_{1}},\dots,\overline{\gamma_{t(n)}}$ correspond to the configurations $\gamma_{1},\dots,\gamma_{t(n)}$, as described below.
We begin by defining the vocabulary $V_{M}$ for $\mathcal{L}_{M}$, which contains all tape symbols $\Gamma$ of $M$ along with a finite set of auxiliary symbols $\overline{\gamma}$ that encode the corresponding configurations $\gamma$. Similar to prior works (Merrill & Sabharwal, 2024), each configuration encoding $\overline{\gamma}$ represents the current state $q$, the symbols at the current head position of $k+2$ tapes (input, output and $k$ work tapes), and the head movement directions $\{0,+1,-1\}$ for each tape. Directions $\{0,+1,-1\}$ denote either staying in place (00), moving left ($-1$), or moving right ($+1$) by a single position. Since the set of states $Q$, the tape alphabet $\Gamma$, and the number of tapes $k$ are all constants, the total number of possible encodings $\overline{\gamma}$ is also constant. Let $\overline{\Gamma}$ denote the set of all possible configuration encodings, i.e., $\overline{\Gamma}=\{\overline{\gamma_{(1)}},\dots,\overline{\gamma_{(l)}}\}$, where $l=|\overline{\Gamma}|$. Given $\overline{\Gamma}$ is finite and enumerable, we can define the rules of the augmented grammar $G_{a}$ accordingly as follows.
@@ 4 CRANE Algorithm
[…]
@@
[…]
@@
[…]
@@ Appendix A Turing Machine Computation
Initially, the machine starts in the initial state $q_{0}\in Q$ with the input tape $c_{0}^{0}$ containing $\boldsymbol{x}$, positioned at index 00, and surrounded by infinite blank symbols ($b$). The head on the input tape is set to $h_{0}^{0}=0$, while all other tapes contain only blank symbols $b$s and have their heads positioned at 00.
@@
where $\delta$ is the transition function. The machine updates each tape $\tau$ by setting $c_{\tau}^{i+1}[h_{\tau}^{i}]=\gamma_{\tau}^{i}$, leaving all other tape cells unchanged. The head position for each tape is updated as $h_{\tau}^{i+1}=h_{\tau}^{i}+d_{\tau}^{i}$. If $q_{i}\in F$, the Turing machine halts and outputs the sequence of tokens on the output tape, starting from the current head position and continuing up to (but not including) the first blank symbol ($b$). A Turing machine can also function as a language recognizer by setting the input alphabet $\Sigma=\{0,1\}$ and interpreting the first output token as either 00 or $1$.
@@ Appendix B Thershold Circuit Class
[…]
@@ Appendix C Proofs
The language $L(G_{c})$ is finite; therefore, for any string $\boldsymbol{s}\in L(G_{c})$, the length satisfies $|\boldsymbol{s}|\leq N$, where $N$ is a constant. Consequently, for any input $\boldsymbol{x}$, the output $\boldsymbol{y}_{G}=\mathcal{L}{}_{G}(\boldsymbol{x})$ has a constant length, i.e., $|\boldsymbol{y}_{G}|\leq N$. The number of autoregressive steps is also bounded by $N$.
@@
($\mathcal{L}{}_{M}(\boldsymbol{x}))=\overline{\gamma_{1}}\cdots\overline{\gamma_{t(n)}}\cdot M(\boldsymbol{x})$. We show that ($\mathcal{L}{}_{M}(\boldsymbol{x})\in L(G_{a})$. $G_{a}\to R_{M}G$. Since, $G$ is output grammar of $M$ then $M(\boldsymbol{x})\in L(G)$. For all $1\leq i\leq t(n)$ $\overline{\gamma_{i}}\in\overline{\Gamma}$. Then, $\overline{\gamma_{1}}\cdots\overline{\gamma_{t(n)}}\in\overline{\Gamma}^{*}\subseteq L(R_{M})$.
Then ($\mathcal{L}{}_{M}(\boldsymbol{x})\in L(G_{a})$ then under constrained decoding the output ($\mathcal{L}{}_{M}(\boldsymbol{x})$ remains unchanged and ($\mathcal{L}{}_{M}(\boldsymbol{x})=\mathcal{L}_{M,{G_{a}}}(\boldsymbol{x})=\boldsymbol{r}\cdot M(\boldsymbol{x})$ where $\boldsymbol{r}=\overline{\gamma_{1}}\cdots\overline{\gamma_{t(n)}}$. ∎

### v4

@@ Abstract
Code generation, symbolic math reasoning, and other tasks require LLMs to produce outputs that are both syntactically and semantically correct. Constrained LLM generation is a promising direction to enforce adherence to formal grammar, but prior works have empirically observed that strict enforcement of formal constraints often diminishes the reasoning capabilities of LLMs. In this work, we first provide a theoretical explanation for why constraining LLM outputs to very restrictive grammars that only allow syntactically valid final answers reduces the reasoning capabilities of the model. Second, we demonstrate that by augmenting the output grammar with carefully designed additional rules, it is always possible to preserve the reasoning capabilities of the LLM while ensuring syntactic and semantic correctness in its outputs. Building on these theoretical insights, we propose a reasoning-augmented constrained decoding algorithm, CRANE, which effectively balances the correctness of constrained generation with the flexibility of unconstrained generation. Experiments on multiple open-source LLMs and benchmarks show that CRANE significantly outperforms both state-of-the-art constrained decoding strategies and standard unconstrained decoding, showing up to 10% points accuracy improvement over baselines on challenging symbolic reasoning benchmarks GSM-symbolic and FOLIO. The code is available at CRANE.
@@ 2.2 Deterministic LLM Decoding
[…]
@@ A Turing machine $M$ with $k$ work tapes and an output tape is a $8$-tuple
where $\Sigma$ is the finite input alphabet, $\Gamma$ is the finite tape alphabet with $\Sigma\subseteq\Gamma$, $b\in\Gamma\setminus\Sigma$ is a special blank symbol, $Q$ is a finite set of states, $q_{0}\in Q$ is the initial state, $\delta:\left(Q\setminus F\right)\times\Gamma^{k+2}\to Q\times\Gamma^{k+1}\times\{0,+1,-1\}^{k+2}$ is the transition function (where 1,1,0-1,1,0 represent moving the tape head left, right, or staying in place, respectively), and $F\subseteq Q$ is the set of halting states.
Let $\Sigma^{*}$ denote the set of all finite strings over the input alphabet $\Sigma$. Given an input string $\boldsymbol{s}\in\Sigma^{*}$, the computation of MMon $s$ is a sequence of configurations starting from the initial configuration. Each configuration $\gamma$ is a tuple containing the current state $q\in Q$, the contents of the input tape, the $k$ work tapes, the output tape, and the current head positions of all $k+2$ tapes. For each configuration, $\gamma_{i}$ ($i\in\mathbb{N}$), the transition function $\delta$ computes the next configuration $\gamma_{i+1}$ based on the current state $q$ and the values on the $k+2$ tapes at the current head positions. It updates the head positions, writes to the output tape (possibly leaving it unchanged if no new symbol is written), and advances to the next configuration. For each $i$, computation of $\gamma_{i+1}$ from $\gamma_{i}$ defines a single step of the Turing machine.
@@ 3.1 Limitation of Constrained Decoding
[…]
@@
For any log-precision LLM $\mathcal{L}{}$ with constant layers there exists a logspace-uniform thershold circuit $Th_{n}$ such that $\mathcal{L}{}_{G_{c}}(\boldsymbol{x})=Th_{n}(\boldsymbol{x})$ holds for all inputs $\boldsymbol{x}$ with size $|\boldsymbol{x}|=n$ and $n\in\mathbb{N}$.
@@ 3.2 Reasoning with Augmented Grammar
[…]
We begin by defining the vocabulary $V_{M}$ for $\mathcal{L}_{M}$, which contains all tape symbols $\Gamma$ of $M$ along with a finite set of auxiliary symbols $\overline{\gamma}$ that encode the corresponding configurations $\gamma$. Similar to prior works (Merrill & Sabharwal, 2024), each configuration encoding $\overline{\gamma}$ represents the current state $q$, the symbols at the current head position of $k+2$ tapes (input, output and $k$ work tapes), and the head movement directions $\{0,+1,-1\}$ for each tape. Directions $\{0,+1,-1\}$ denote either staying in place (0), moving left (1-1), or moving right (+1+1) by a single position. Since the set of states $Q$, the tape alphabet $\Gamma$, and the number of tapes $k$ are all constants, the total number of possible encodings $\overline{\gamma}$ is also constant. Let $\overline{\Gamma}$ denote the set of all possible configuration encodings, i.e., $\overline{\Gamma}=\{\overline{\gamma_{(1)}},\dots,\overline{\gamma_{(l)}}\}$, where $l=|\overline{\Gamma}|$. Given $\overline{\Gamma}$ is finite and enumerable, we can define the rules of the augmented grammar $G_{a}$ accordingly as follows.
@@ 4 CRANE Algorithm
8:  $\text{currGen}\leftarrow\text{detokenize}(\text{tokens}[\texttt{pointer}:])$
9:  if $\text{$S_{1}$}\in\texttt{currGen}$ then
10:   $\texttt{isConstrained}\leftarrow\text{True}$
@@
14:   $\text{constrained}\leftarrow\text{extractConstrained}(\texttt{currGen})$
@@
24:     $\texttt{pointer}\leftarrow\text{len}(\text{tokens})$
25:return $\text{detokenize}(\text{tokens})$
@@ Appendix A Turing Machine Computation
Initially, the machine starts in the initial state $q_{0}\in Q$ with the input tape $c_{0}^{0}$ containing $\boldsymbol{x}$, positioned at index 0, and surrounded by infinite blank symbols ($b$). The head on the input tape is set to $h_{0}^{0}=0$, while all other tapes contain only blank symbols $b$s and have their heads positioned at 0.
@@
where $\delta$ is the transition function. The machine updates each tape $\tau$ by setting $c_{\tau}^{i+1}[h_{\tau}^{i}]=\gamma_{\tau}^{i}$, leaving all other tape cells unchanged. The head position for each tape is updated as $h_{\tau}^{i+1}=h_{\tau}^{i}+d_{\tau}^{i}$. If $q_{i}\in F$, the Turing machine halts and outputs the sequence of tokens on the output tape, starting from the current head position and continuing up to (but not including) the first blank symbol ($b$). A Turing machine can also function as a language recognizer by setting the input alphabet $\Sigma=\{0,1\}$ and interpreting the first output token as either 0 or $1$.
@@ Appendix B Thershold Circuit Class
$TC^{0}$ is a class of computational problems that can be recognized by constant-depth, polynomial-size circuits composed of threshold gates. A threshold gate, such as $\theta_{\leq k}$, outputs 1 if the sum of its input bits is at most $k$, while $\theta_{\geq k}$ outputs 1 if the sum is at least $k$. These circuits also include standard logic gates like $\wedge$, $\vee$, and $\neg$ as special cases of threshold functions. Since $TC^{0}$ circuits can simulate $AC^{0}$ circuits ( a polysize, constant-depth $\{\wedge,\vee,\neg\}$-circuit family), they are at least as powerful as $AC^{0}$ in the computational hierarchy. The circuit families we have defined above are non-uniform, meaning that there is no requirement for the circuits processing different input sizes to be related in any way. In degenerate cases, non-uniform circuit families can solve undecidable problems making them an unrealizable model of computation (Arora & Barak, 2009). Intuitively, a uniform circuit family requires that the circuits for different input sizes must be ”somewhat similar” to each other. This concept is formalized by stating that there exists a resource-constrained Turing machine that, given the input $1^{n}$, can generate a serialization of the corresponding circuit $C_{n}$ for that input size. Specifically, a logspace uniform $TC^{0}$ family can be constructed by a logspace-bounded Turing machine from the string $1^{n}$.
@@ Appendix C Proofs
[…]
@@
$\mathcal{L}{}_{M}(\boldsymbol{x}))=\overline{\gamma_{1}}\cdots\overline{\gamma_{t(n)}}\cdot M(\boldsymbol{x})$. We show that $\mathcal{L}{}_{M}(\boldsymbol{x})\in L(G_{a})$. $G_{a}\to R_{M}G$. Since, $G$ is output grammar of $M$ then $M(\boldsymbol{x})\in L(G)$. For all $1\leq i\leq t(n)$ $\overline{\gamma_{i}}\in\overline{\Gamma}$. Then, $\overline{\gamma_{1}}\cdots\overline{\gamma_{t(n)}}\in\overline{\Gamma}^{*}\subseteq L(R_{M})$.
[…]